OpenStudy (anonymous):

A caterer charges the same price for each wrap ordered. Each bottle of water is also the same price. The cost of buying 50 wraps and 60 bottles of water is $445. The cost of buying 60 wraps and 40 bottles of water is $470. What is the price of each? a. Each wrap costs $2.00, and each bottle of water costs $6.50. b. Each wrap costs $4.70, and each bottle of water costs $4.70. c. Each wrap costs $5.30, and each bottle of water costs $3.00. d. Each wrap costs $6.50, and each bottle of water costs $2.00.

OpenStudy (anonymous):

@RyleeC I didnt pass that one , so I have a couple more .. if thats ok

OpenStudy (anonymous):

50w+60x=445 60w+40x=470

OpenStudy (anonymous):

w=8.9-1.2x

OpenStudy (anonymous):

60(8.9-1.2x)+40x=470 solve for x

OpenStudy (anonymous):

$2

OpenStudy (anonymous):

So A ?

OpenStudy (anonymous):

Yep.